Cellular respiration is the process that allows your body to harvest a huge amount of energy from a single glucose. In fact, it's so efficient that you get 30 ATP for every 1 glucose molecule you eat. This exercise is going to break down exactly where each ATP comes from. Instructions: Walk the audience through every step that is involved in generating an ATP, NADH or FADH2, include the names, structures and enzyme for each step. From there, look up the conversion rates from NADH and FADH2 to ATP and calculate exactly how 30 ATP are formed. Also include any step which uses an ATP (this will count against the total ATP formed).
Answer #1

Ans)

Note :- values inside brackets represent ATP synthesised as per new P:O ratios of 2.5 for NADH and 1.5 for FADH2. I have taken1 NADH =3 ATP and 1 FADH= 2 ATP, and 1GTP = 1ATP for ease of calculation), also * is for multiplication.

Glyceraldehyde - 3 phosphate dehydrogenase. =2 NADH, so 2*3= 6 (5)ATP (1 NADH = 3ATP).

Phosphoglycsrate kinase = 2 ATP

Pyruvate kinase = 2ATP

Two ATP consumed in the reaction catalysed by hwxokinase and phospgofructokinase =(-2)

Net ATP synthesis in glycolysis is 8​​​​​​(7).

Pyruvate dehydrogenase = 2 * 3ATP =6 ATP. ( 2 because 2 molecules of pyruvate is formed from glucose and conversion of each pyruvate gives off 1 NADH)

Krebs cycle

* isocitrate dehydrogenase = 2NADH so 6(5) ATP.

*alpha keto glutarate dehydrogenase =2 NADH so 6 (5) ATP.

* succinate thiokinase = 2 GTP so 2 ATP

* succinate dehydrogenase = 2 FADH2 so 2*2 = 4(3)ATP .

*malate dehydrogenase = 2 NADH so 2*3ATP =6(5) ATP.

So total ATP per mole oh glucose under aerobic condition is 38 (32) ATP.